# Towards Designing a Question-Answering Chatbot for Online News: Understanding Perspectives and Questions

### 17 Dec 2023 v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/1122445.1122456conference: Proc. ACM Hum.-Comput. Interact.; June 03–05, 2018; Woodstock, NYbooktitle: Proc. ACM Hum.-Comput. Interactprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Journalists and reporters carefully integrate information and interpretation in a news article to inform readers about a topic, event, or phenomenon in a balanced and objective manner (Mencher and Shilton, 1997; Campbell, 1997; Maras, 2013; Starkey, 2017). Nonetheless, prior research shows that oftentimes readers are left with questions for journalists/authors after reading news (Castellano Parra et al., 2020; Weber, 2014). Readers use various means such as news comment sections, email, and social media direct messages to ask their questions (Diakopoulos and Naaman, 2011; Martin and Murrell, 2021). However, due to many reasons such as the abundance of questions, lack of time, and absence of direct incentives, journalists cannot respond to the queries (Cohen, 2019; Monteiro et al., 2016). On the other hand, both journalists and newsrooms have an interest in engaging with the audience/readers, to learn what aspects of their news raised curiosity and questions in readers’ minds (Meier et al., 2018; Nelson, 2021), and to understand if readers are questioning newsroom’s journalistic integrity (Masullo et al., 2017). This knowledge is critical for learning what attracts readers to news content, targeting the right audience for news recommendations, building trust with the audience and so on (Li et al., 2010; Wu et al., 2020). According to Meier et al. (Meier et al., 2018),- “a paradigm shift away from a ‘lecturing’ approach to a ‘dialogue’ approach is a key factor for journalism in a post-truth era”. So, there is a clear interest from both readers’ and newsrooms’ perspectives to engage with each other in a dialogue fashion through question-answering (Meier et al., 2018; Peña Fernández et al., 2021).
@@
With the rise of Large Language Models (LLMs) (Brown et al., 2020), like any other industry, news media industry is also investigating how Artificial Intelligence (AI) and LLMs can be integrated with the news production, distribution, audience engagement, and other related processes to improve efficiency, increase trust and accountability, and open new possibilities (Veglis and Maniou, 2019b; Hong and Oh, 2020a; Veglis and Maniou, 2019a). For instance, BBC has started experimenting with how we can leverage bot technologies to reach new audiences on messaging platforms and social media (BBC News Labs, 2023a, b). They have developed a prototype of an in-article chatbot to help less engaged audiences understand big news stories (bbc, 2018). Other media outlets have started experimenting with Meta (previously known as Facebook) chatbots (Zhang et al., 2021a). While these experiments are ongoing individually in different newsrooms, facilitating question-answering (QA), one of the primary modes of interaction and dialogue in chatbots (Liao et al., 2018), remains a challenge for newsrooms (Trischler et al., 2017; The Verge, 2023). While researchers are working towards mitigating technological limitations such as the lack of truthfulness and existence of biased and formulaic text in LLM responses (Weidinger et al., 2021; Epstein et al., 2023), we argue that it is also important to take a systematic approach to understand how journalists and readers want to use a chatbot for question-answering and what roles they deem fit for such a chatbot. For instance, are journalists comfortable with a chatbot answering readers’ questions about their written news article? What types of questions readers expect the chatbot to answer? And what types of questions readers expect the authors to answer? Without answering these questions, we may end up designing QA chatbots that are detached from readers’ and newsrooms’ needs.
@@
This paper takes a formative step towards designing LLM-powered chatbots that will be able to answer open-ended questions from readers, reduce journalists burden to answer a large number of queries, and at the same time elevate communication between readers and journalist. We call such chatbots as QA chatbots. For the rest of the paper, we use the term QA chatbots and chatbots interchangeably to refer to the chatbots that are able to answer questions from news readers, unless specified otherwise. While we only focus on question-answering in the context of LLMs, in a real-world setting, like many LLM-powered chatbots (e.g., ChatGPT), a QA chatbot (or simply a chatbot) will likely have other forms of interactions (e.g., recommending news). To achieve our research agenda, we conducted two studies in this paper. First, we interviewed six professional journalists to understand how journalists currently answer questions from readers and, more broadly, interact with readers. The interviews revealed that most journalists now obtain feedback and answer questions from readers through one-to-one conversations over email, direct messages, or in-person meetups, except for a few member gatherings organized by the news outlets. All participants were enthusiastic about using a QA chatbot as a mediator between themselves and readers. They want the chatbot to answer factual and redundant questions from readers while want the chatbot to direct questions seeking subjective interpretation towards them.
@@
We subsequently conducted a between-subject experiment on Amazon Mechanical Turk (MTurk) with 124 participants with two goals: (1) understanding to what extent readers’ perspectives and questioning patterns match with how journalists see the functionalities of a QA chatbot; and (2) how would a QA chatbot modulate readers’ questioning patterns. The second goal is motivated from prior research that shows that chatbots can significantly modulate user behaviour (Luger and Sellen, 2016; Chaves and Gerosa, 2021; Liao et al., 2016). In the study, participants were instructed to read three articles from three different domains (health, politics, and environment) and ask questions to the authors of the articles or a chatbot. We analyzed their responses using a grounded theory approach (Glaser and Strauss, 2017) and a set of quantitative measures. Our findings suggest that readers mostly asked questions on two broad categories: Information, questions that seek short-form factual and long-form details; and Interpretation, questions that seek explanation and opinion. We found that readers’ questioning patterns significantly differ depending on the receiver - authors vs. chatbot. Participants asked questions seeking facts more frequently to the chatbot than to the authors. In contrast, they asked questions seeking explanation and opinion with multiple facets more frequently to authors than to the chatbot.
@@
These results indicate that readers’ perspectives about the role of QA chatbots match with journalists’ view to some extent. Readers too predominantly want the chatbot to answer factual questions while seeking to engage in subjective conversations with journalists. Despite this alignment, we found evidence that a significant number of readers do not share the view that chatbots should only answer factual questions. For instance, while not as frequently as to the authors, readers did ask a considerable amount of questions seeking subjective interpretation to the chatbot. Similarly, a considerable number of readers wanted authors to answer factual as well as subjective questions. Finally, we found that chatbot can modulate readers’ behavior significantly which researchers and news organizations should be aware of. For example, articles having lower perceived quality motivated readers to be critical and question journalistic integrity when the authors were present in the loop. However, when the chatbot was present, readers did not engage on the same level, even when the perceived quality of the news article was low.
@@
Overall, these results indicate that researchers and news organizations need to understand journalists’ and readers’ perspectives and expectations thoroughly before determining the functionalities of a QA chatbot. The results also highlight the need to devise policies around chatbots and communicate the policies to journalists and readers effectively. We conclude this paper by proposing a framework for devising policy for designing QA chatbots for online news. In summary, our contributions are as follows: (1) An interview study with six journalists to understand their current practices and challenges for answering questions from readers and how they perceive the role of an intelligent chatbot in this regard; (2) An online experiment with 124 participants to understand what questions readers typically ask about a news article and how the questions change if a chatbot is present at the receiving end, instead of the authors; and (3) A design framework for developing a chatbot policy, informed by the interviews, online experiment, and literature review.
@@ 2.1. Audience Engagement and Chatbots
Social media has drastically changed audience engagement and news consumption paradigm (e.g., Twitter) (Xia et al., 2020; Wu et al., 2019). About half of the U.S. population now access news from social media (Paw Research Center, 2021). Thus, news media are no longer the authoritative sources of news (Boczkowski et al., 2018). With this shift in dynamics, news organizations are exploring different ways to adapt to social media environments and engage audience (Hong and Oh, 2020b). One of the promising ways is designing chatbots in social media (Hong and Oh, 2020b; Lokot and Diakopoulos, 2016). For example, several large international news organizations such as ABC News, NBC News, and BBC News have chatbots in Facebook messenger (Zhang et al., 2021b). These chatbots, including others in this area (Lokot and Diakopoulos, 2016; Maniou and Veglis, 2020; Zhang et al., 2021b), largely focus on supporting fact-checking, news dissemination, and recommendation.
@@ 3.2. Method
We conducted two interviews in person and others over Zoom. One author of this paper conducted the interviews while another author took notes. Each interview was divided into three parts. First, after gathering informed consent, we asked participants to describe how they currently interact with readers, answer questions from readers, and manage conversations with readers. After that, we asked participants about any challenges they faced in the process. Finally, participants brainstormed with the authors about how a chatbot can enrich journalist-reader communication. Each session lasted around 1 hour. The semi-structured questionnaire is available in an anonymized OSF repository.
@@ 4. Study with Audience/Readers
While the perspective of journalists is clear from the interviews, to design an effective QA chatbot, we still need to know how readers, the other significant stakeholder, perceive chatbots. More importantly, since journalists want to use the chatbot as a mediator, we wanted to know to what extent readers’ questioning patterns match with this view. We determined that conducting a study to understand how readers want to route questions between the authors and chatbot would be essential to achieve this. We also wanted to understand how a chatbot might influence the questioning patterns from readers. Prior research suggests chatbots can significantly modulate user behaviour (Luger and Sellen, 2016; Chaves and Gerosa, 2021; Liao et al., 2016). Thus, we seek to answer the following two research questions:
@@
To answer the RQs, we decided to conduct a crowdsourced experiment. While a contextual interview study with readers could also be useful here, we decided that a crowdsourced experiment is better suited here since it would enable us to understand the perspectives of a larger population, engage readers in the task of concern (asking questions), and help us answer the RQs. This section describes the study design and protocol. All study materials, including news articles and the source code for the study interface, are available in the anonymized OSF repository. The study took place in June 2023.
@@ 4.1. Study Conditions
We also considered the possibility of any hidden confounds. For instance, there are two differences between A and C: human vs AI and authors vs non-authors. We did not study these facets separately as our goal was to understand how readers direct questions towards authors and chatbot, not how the questions differ between human vs AI or authors vs non-authors. In other words, we are interested in the complete constructs embodied by authors and chatbot, not their individual components. Additionally, assuming AI as authors could lead to ethical problems associated with anthropomorphizing AI (Deshpande et al., 2023); AI models are not persons and thus cannot be authors in the true sense, and there are legal, safety, security, trust, and reliability concerns in such relationships (Deshpande et al., 2023; Epstein et al., 2023).
@@ 4.2. Participants
The details of the participants are provided in Figure 7. MTurk has several embedded functionalities to control the quality of participants. We required participants to be located in the United States (US), have at least 50 approved tasks, and be Master workers, a qualification assigned to workers who have demonstrated a high rate of success in completing a wide range of tasks. On average, it took 29 minutes for participants to complete the study. Considering the minimum wage standards in our US State, we compensated each participant \$8 for their time.
@@ 4.3. News Articles and Collection Method
By employing this systematic approach to article selection, we ensured that our study incorporated a wide variety of perspectives and minimized potential biases in the articles analyzed. The final counts for the source news outlets are– Reuters: 1, BBC News: 2, NBC: 1, NPR: 1, AP News: 1, CNN: 1, NY Times: 2, Washington Times: 1, NY Post: 2, Fox News: 2, Daily Mail: 1. To avoid potential bias towards a news organization, we removed all identifiable information such as Author’s name from the articles and replaced reference to the news organizations name with generic names. For each participant in the study, we randomly chose one article from each domain. Figure 8 shows the number of times the articles were read by the participants.
@@ 4.4. Study Interface
4.5. Protocol
  1. (1)
  2. (2)
    1. (a)
      Read the article: Participants read the article in a PDF reader.
    2. (b)
    3. (c)
    4. (d)
  3. (3)
4.6. Quality Control
4.7. Analysis
@@ 5. Results from the Study with Audience
We collected 752 questions (3 articles * 2 questions * 124 participants = 744 + 8 extra) from MTurk. This section presents the results of our study. We structure the results around our research questions (RQs).
@@ 5.1. A Taxonomy of Questions (RQ1)
Figure 2. Number of questions asked to the authors and chatbot.
This figure shows a bar chart with multi-colored bars.
Figure 2. Number of questions asked to the authors and chatbot.

### 24 Mar 2024 v2

# Towards Designing a Question-Answering Chatbot for Online News: Understanding Questions and Perspectives

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642007isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Journalists and reporters carefully integrate information and interpretation in a news article to inform readers about a topic, event, or phenomenon in a balanced and objective manner (Mencher and Shilton, 1997; Campbell, 1997; Maras, 2013; Starkey, 2017). Nonetheless, prior research shows that oftentimes readers are left with questions for journalists/authors after reading news (Castellano Parra et al., 2020; Weber, 2014). Readers use various means such as news comment sections, email, and social media direct messages to ask their questions (Diakopoulos and Naaman, 2011; Martin and Murrell, 2021). However, due to many reasons such as the abundance of questions, lack of time, and absence of direct incentives, journalists cannot respond to the queries (Cohen, 2019; Monteiro et al., 2016). On the other hand, both journalists and newsrooms have an interest in engaging with the audience/readers, to learn what aspects of their news raised curiosity and questions in readers’ minds (Meier et al., 2018; Nelson, 2021), and to understand if readers are questioning newsroom’s journalistic integrity (Masullo et al., 2017). This knowledge is critical for learning what attracts readers to news content, targeting the right audience for news recommendations, building trust with the audience and so on (Li et al., 2010; Wu et al., 2020). According to Meier et al. (Meier et al., 2018),- “a paradigm shift away from a ‘lecturing’ approach to a ‘dialogue’ approach is a key factor for journalism in a post-truth era”. Thus, there is a clear interest from both readers’ and newsrooms’ perspectives to engage with each other in a dialogue fashion through question-answering (Meier et al., 2018; Peña Fernández et al., 2021).
@@
With the rise of Large Language Models (LLMs) (Brown et al., 2020), like any other industry, news media industry is also investigating how Artificial Intelligence (AI) and LLMs can be integrated with the news production, distribution, audience engagement, and other related processes to improve efficiency, increase trust and accountability, and open new possibilities (Veglis and Maniou, 2019b; Hong and Oh, 2020a; Veglis and Maniou, 2019a). For instance, BBC has started experimenting with how we can leverage bot technologies to reach new audiences on messaging platforms and social media (BBC News Labs, 2023a, b). They have developed a prototype of an in-article chatbot to help less engaged audiences understand big news stories (Labs, 2018). Other media outlets have started experimenting with Meta (previously known as Facebook) chatbots (Zhang et al., 2021). While these experiments are ongoing individually in different newsrooms, facilitating question-answering (QA), one of the primary modes of interaction and dialogue in chatbots (Liao et al., 2018), remains a challenge for newsrooms (Trischler et al., 2017; The Verge, 2023). While researchers are working towards mitigating technological limitations such as the lack of truthfulness and existence of biased and formulaic text in LLM responses (Weidinger et al., 2021; Epstein et al., 2023), we argue that it is also important to take a systematic approach to understand how journalists and readers want to use a chatbot for question-answering and what roles they deem fit for such a chatbot. For instance, are journalists comfortable with a chatbot answering readers’ questions about their written news article? What types of questions do readers expect the chatbot to answer? And what types of questions do readers expect the authors to answer? Without answering these questions, we may end up designing QA chatbots that are detached from readers’ and newsrooms’ needs.
@@
This paper takes a formative step towards designing LLM-powered chatbots that will be able to answer open-ended questions from readers, reduce journalists’ burden to answer a large number of queries, and at the same time elevate communication between readers and journalists. We call such chatbots as QA chatbots. For the rest of the paper, we use the term QA chatbots and chatbots interchangeably to refer to the chatbots that are able to answer questions from news readers, unless specified otherwise. While we only focus on question-answering in the context of LLMs, in a real-world setting, like many LLM-powered chatbots (e.g., ChatGPT), a QA chatbot (or simply a chatbot) will likely have other forms of interactions (e.g., recommending news). To achieve our research agenda, we conducted two studies in this paper. First, we interviewed six professional journalists to understand how journalists currently answer questions from readers and, more broadly, interact with readers. The interviews revealed that most journalists now obtain feedback and answer questions from readers through one-to-one conversations over email, direct messages, or in-person meetups, except for a few member gatherings organized by the news outlets. All participants were enthusiastic about using a QA chatbot as a mediator between themselves and readers. They want the chatbot to answer factual and redundant questions from readers while want the chatbot to direct questions seeking subjective interpretation towards them.
@@
We subsequently conducted a between-subject experiment on Amazon Mechanical Turk (MTurk) with 124 participants with two goals: (1) understanding to what extent readers’ perspectives and questioning patterns match with how journalists see the functionalities of a QA chatbot; and (2) understanding how a QA chatbot might modulate readers’ questioning patterns. The second goal is motivated by prior research that shows that chatbots can significantly modulate user behaviour (Luger and Sellen, 2016; Chaves and Gerosa, 2021; Liao et al., 2016). In the study, participants were instructed to read three articles from three different domains (health, politics, and environment) and ask questions to the authors of the articles or a chatbot. We analyzed their responses using a grounded theory approach (Glaser and Strauss, 2017) and a set of quantitative measures. Our findings suggest that readers mostly asked questions on two broad categories: Information, questions that seek short-form factual and long-form details; and Interpretation, questions that seek explanation and opinion. We found that readers’ questioning patterns significantly differ depending on the receiver - authors vs. chatbot. Participants asked questions seeking facts more frequently to the chatbot than to the authors. In contrast, they asked questions seeking explanation and opinion with multiple facets more frequently to authors than to the chatbot.
@@
These results indicate that readers’ perspectives about the role of QA chatbots match with journalists’ views to some extent. Readers too predominantly want the chatbot to answer factual questions while seeking to engage in subjective conversations with journalists. Despite this alignment, we found evidence that a significant number of readers do not share the view that chatbots should only answer factual questions. For instance, while not as frequently as to the authors, readers did ask a considerable amount of questions seeking subjective interpretation to the chatbot. Similarly, a considerable number of readers wanted authors to answer factual as well as subjective questions. Finally, we found that chatbots can modulate readers’ behavior significantly which researchers and news organizations should be aware of. For example, articles having lower perceived quality motivated readers to be critical and question journalistic integrity when the authors were present in the loop. However, when the chatbot was present, readers did not engage on the same level, even when the perceived quality of the news article was low.
@@
Overall, these results indicate that researchers and news organizations need to understand journalists’ and readers’ perspectives and expectations thoroughly before determining the functionalities of a QA chatbot. The results also highlight the need to devise policies around chatbots and communicate the policies to journalists and readers effectively. We conclude this paper by proposing a framework for devising a policy for designing QA chatbots for online news. In summary, our contributions are as follows: (1) An interview study with six journalists to understand their current practices and challenges for answering questions from readers and how they perceive the role of an intelligent chatbot in this regard; (2) An online experiment with 124 participants to understand what questions readers typically ask about a news article and how the questions change if a chatbot is present at the receiving end, instead of the authors; and (3) A design framework for developing a chatbot policy, informed by the interviews, online experiment, and literature review.
@@ 2.1. Audience Engagement and Chatbots
Social media has drastically changed audience engagement and news consumption paradigm (e.g., Twitter) (Xia et al., 2020; Wu et al., 2019). About half of the U.S. population now access news from social media (Paw Research Center, 2021). Thus, news media are no longer the authoritative sources of news (Boczkowski et al., 2018). With this shift in dynamics, news organizations are exploring different ways to adapt to social media environments and engage audience (Hong and Oh, 2020b). One of the promising ways is designing chatbots in social media (Hong and Oh, 2020b; Lokot and Diakopoulos, 2016). For example, several large international news organizations such as ABC News, NBC News, and BBC News have chatbots in Facebook messenger (Zhang et al., 2021). These chatbots, including others in this area (Lokot and Diakopoulos, 2016; Maniou and Veglis, 2020; Zhang et al., 2021), largely focus on supporting fact-checking, news dissemination, and recommendation.
@@ 3.2. Method
We conducted two interviews in person and others over Zoom. One author of this paper conducted the interviews while another author took notes. Each interview was divided into three parts. First, after gathering informed consent, we asked participants to describe how they currently interact with readers, answer questions from readers, and manage conversations with readers. After that, we asked participants about any challenges they faced in the process. Finally, participants brainstormed with the authors about how a chatbot can enrich journalist-reader communication. Each session lasted around 1 hour. The semi-structured questionnaire is available in our OSF repository.
@@ 4. Study with Audience/Readers
While the perspective of journalists is clear from the interviews, to design an effective QA chatbot, we still need to know how readers, the other significant stakeholder, perceive chatbots. More importantly, since journalists want to use the chatbot as a mediator, we wanted to know to what extent readers’ questioning patterns match with this view. We determined that conducting a study to understand how readers want to route questions between the authors and the chatbot would be essential to achieve this. We also wanted to understand how a chatbot might influence the questioning patterns of readers. Prior research suggests chatbots can significantly modulate user behaviour (Luger and Sellen, 2016; Chaves and Gerosa, 2021; Liao et al., 2016). Thus, we seek to answer the following two research questions:
@@
To answer the RQs, we decided to conduct a crowdsourced experiment. While an interview study with readers could also be useful here, we decided that a crowdsourced experiment is better suited here since it would enable us to understand the perspectives of a larger population, engage readers in the task of concern (asking questions), and help us answer the RQs. This section describes the study design and protocol. All study materials, including news articles and the source code for the study interface, are available in our OSF repository. The study took place in June 2023.
@@ 4.2. Participants
The details of the participants are provided in Figure 7 (Appendix A). MTurk has several embedded functionalities to control the quality of participants. We required participants to be located in the United States (US), have at least 50 approved tasks, and be Master workers, a qualification assigned to workers who have demonstrated a high rate of success in completing a wide range of tasks. On average, it took 29 minutes for participants to complete the study. Considering the minimum wage standards in our US State, we compensated each participant \$8 for their time.
@@ 4.3. News Articles and Collection Method
By employing this systematic approach to article selection, we ensured that our study incorporated a wide variety of perspectives and minimized potential biases in the articles analyzed. The final counts for the source news outlets are– Reuters: 1, BBC News: 2, NBC: 1, NPR: 1, AP News: 1, CNN: 1, NY Times: 2, Washington Times: 1, NY Post: 2, Fox News: 2, Daily Mail: 1. To avoid potential bias towards a news organization, we removed all identifiable information such as Author’s name from the articles and replaced reference to the news organizations name with generic names. For each participant in the study, we randomly chose one article from each domain. Figure 8 (Appendix A) shows the number of times the articles were read by the participants.
@@ 4.4. Study Interface
4.5. Protocol
  1. (1)
  2. (2)
    1. (a)
      Read the article: Participants read the article in a PDF reader.
    2. (b)
    3. (c)
    4. (d)
  3. (3)
4.6. Quality Control
4.7. Analysis
Figure 2. Number of questions asked to the authors and chatbot.
This figure shows a bar chart with multi-colored bars.
Figure 2. Number of questions asked to the authors and chatbot.
@@ 5. Results from the Study with Audience
We collected 752 questions (3 articles * 2 questions * 124 participants = 744 + 8 extra) from MTurk. This section presents the results of our study. We structure the results around our RQs.
@@ 5.1. A Taxonomy of Questions (RQ1)
This figure shows multiple visual elements and mapping between them.
@@ 5.2.2. Complexity
Table 3. Results of the SEM model. Each cell represents the coefficient for the respective relation. Same coefficient values between the author and chatbot conditions indicate globally constrained paths in the SEM model (i.e., no differences between the groups) based on the multigroup analysis. p-value significance: * p ¡ 0.05, ** p ¡ 0.01, *** p ¡ 0.001.
Authors Chatbot
Fact
-0.13*
0.04
-0.02
-0.03
-0.14
0.04
-0.02
-0.03
Details
0.42***
-0.07
-0.09
0.15
0.35**
-0.07
-0.09
0.15
Evidence
-0.40***
0.05
0.03
-0.09
0.10
0.05
0.03
-0.096
Closed-ended
0.003
-0.12
0.16
0.25
0.12*
-0.12
0.16
0.25
Opinion
0.22*
0.01
0.05
-0.07
0.04
0.01
0.05
-0.07
Explanation
-0.27**
0.04
0.08
-0.08
0.06
0.04
0.08
-0.08
Journalistic Integrity
-0.54***
0.07
0.06
-0.13
-0.11
0.07
0.06
-0.13
Bloom’s Taxonomy
0.06
0.20
-0.03
-0.17
0.14
0.20
-0.03
-0.17
Length
14.70
24.22
3.60
30.04
11.23
24.22
3.60
30.04
Sentiment
-0.03
0.007
-0.03
-0.001
-0.05
0.007
-0.03
-0.001
Table 3. Results of the SEM model. Each cell represents the coefficient for the respective relation. Same coefficient values between the author and chatbot conditions indicate globally constrained paths in the SEM model (i.e., no differences between the groups) based on the multigroup analysis. p-value significance: * p ¡ 0.05, ** p ¡ 0.01, *** p ¡ 0.001.